OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -625 and a2 = 125

OpenStudy (cwrw238):

first find the common ratio of the sequence - can you do that?

rishavraj (rishavraj):

find the common ratio first of all...nd then find the seventh term

OpenStudy (anonymous):

is it -500

rishavraj (rishavraj):

no its not tht

OpenStudy (cwrw238):

the seventh term = a1*r^6 where r = common ratio

OpenStudy (cwrw238):

r = a2 / a1

OpenStudy (anonymous):

what is the common difference

OpenStudy (anonymous):

-5

OpenStudy (cwrw238):

common ratio = 125 / -625

OpenStudy (anonymous):

-0.2

OpenStudy (cwrw238):

coomon differnce is a property of an arithmetic sequence geometric sequence has a common ratio

rishavraj (rishavraj):

common ratio means a2/a1

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

is the answer 0.04

OpenStudy (cwrw238):

-0/2 is correct for r 7th term = -625 * (-0.2)^6

rishavraj (rishavraj):

125 / -625 = -1/5

OpenStudy (anonymous):

so -0.04

OpenStudy (cwrw238):

* sorry that' s r = -0.2 or -1/5

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

thank you guys!

rishavraj (rishavraj):

@Johnngorskii -0.04 is correct